# Weighted Particle-Based Optimization for Efficient Generalized Posterior Calibration

### v1

@@ Abstract
In the realm of statistical learning, the increasing volume of accessible data and increasing model complexity necessitate robust methodologies. This paper explores two branches of robust Bayesian methods in response to this trend. The first is generalized Bayesian inference, which introduces a learning rate parameter to enhance robustness against model misspecifications. The second is Gibbs posterior inference, which formulates inferential problems using generic loss functions rather than probabilistic models. In such approaches, it is necessary to calibrate the spread of the posterior distribution by selecting a learning rate parameter. The study aims to enhance the generalized posterior calibration (GPC) algorithm proposed by Syring and Martin (2019) [Biometrika, Volume 106, Issue 2, pp. 479-486]. Their algorithm chooses the learning rate to achieve the nominal frequentist coverage probability, but it is computationally intensive because it requires repeated posterior simulations for bootstrap samples. We propose a more efficient version of the GPC inspired by sequential Monte Carlo (SMC) samplers. A target distribution with a different learning rate is evaluated without posterior simulation as in the reweighting step in SMC sampling. Thus, the proposed algorithm can reach the desired value within a few iterations. This improvement substantially reduces the computational cost of the GPC. Its efficacy is demonstrated through synthetic and real data applications.
1 Introduction
@@
$\pi\left(\boldsymbol{\theta}\right)\propto p\left(\mathcal{D}|\boldsymbol{\theta}\right)p\left(\boldsymbol{\theta}\right).$
In contrast, generalized Bayesian inference introduces a generalized posterior by incorporating a learning rate $\eta\left(>0\right)$ (also termed a scaling parameter) alongside the likelihood, as follows:
@@
This paper proposes an improvement in the GPC that requires fewer of iterations. This algorithm is inspired by sequential Monte Carlo (SMC) samplers [18], exploiting the similarity between the learning rate in the GPC and an inverse temperature in SMC sampling. The target distribution with the current value of $\eta$ is approximated using weighted particles, as in SMC sampling. The particle approximation of the target distribution with a new value of $\eta$ is obtained without posterior simulation as in the reweighting step in SMC sampling. At least in theory, a desired value of $\eta$ is reached in a single step, although few iterations are required in practice due to the particle degeneracy problem.
A recent paper [19] proposed an algorithm that replaces an MCMC sampler with an SMC sampler. The algorithm proposed in this paper is different, in that it uses an MCMC sampler for posterior simulation and improves on the optimization procedure itself. SMC samplers are only applicable to relatively low-dimensional cases (say, with a sampling dimension less than 10). In contrast, there are MCMC samplers that are feasible for high-dimensional cases.
@@
2 Background
2.1 Generalized posterior calibration
@@
where $\mathbb{I}\left\{\cdot\right\}$ denotes the indicator function. $\eta$ is selected by solving $\hat{c}_{\alpha}\left(\eta|\mathbb{P}_{N}\right)=1-\alpha$ via a stochastic approximation [20]. At the $s$th iteration, a single step of stochastic approximation recursion is applied:
@@
where $\left\{\varsigma_{l}\right\}$ is a nonincreasing sequence such that $\sum_{l}\varsigma_{l}=\infty$ and $\sum_{l}\varsigma_{l}^{2}<\infty$. While [16] specified $\varsigma_{s}=s^{-0.51}$, we adopt a variant of Keston’s [21] rule (Figure 2), $\varsigma_{s}\left(l\right)=l^{-0.51}$, where $l$ increases by one only when there is a directional change in the trajectory of $\eta_{s}$ and $\hat{c}_{\alpha}\left(\eta_{s}|\mathbb{P}_{N}\right)<1$. This adaptation significantly reduces the convergence time.
@@
2.2 Sequential Monte Carlo sampler
The following describles sequential Monte Carlo (SMC) samplers [18]. These algorithms repeatedly utilize importance sampling to generate sets of weighted particles $\left\{w^{\left[m\right]},\boldsymbol{\theta}^{\left[m\right]}\right\}_{m=1}^{M}$ that approximate a sequence of synthetic intermediate distributions $\left\{\pi_{t}\right\}_{t=0}^{T}$. Initially, the distribution $\pi_{0}$ corresponds to the prior, $\pi_{0}\left(\boldsymbol{\theta}\right)=p\left(\boldsymbol{\theta}\right)$, while the terminal distribution $\pi_{T}$ is the target distribution, i.e., the posterior, $\pi_{T}\left(\boldsymbol{\theta}\right)=p\left(\mathcal{D}|\boldsymbol{\theta}\right)p\left(\boldsymbol{\theta}\right)$. Each intermediate distribution is defined as a likelihood-tempered posterior:
@@ 2.2 Sequential Monte Carlo sampler
where $\left\{\phi_{t}\right\}_{t=1}^{T}$ is an increasing sequence with $\phi_{0}=0$ and $\phi_{T}=1$. The parameter $\phi_{t}$ can be interpreted as the inverse temperature. We denote a sequence of auxiliary distributions as $\left\{\tilde{\pi}_{t}\right\}_{t=0}^{T}$ with
$\tilde{\pi}_{t}\left(\boldsymbol{\theta}_{0:t}\right)=\pi_{t}\left(\boldsymbol{\theta}_{t}\right)\prod_{s=0}^{t-1}\mathcal{L}_{s}\left(\boldsymbol{\theta}_{s+1},\boldsymbol{\theta}_{s}\right),$
where $\mathcal{L}_{t}\left(\cdot,\cdot\right)$ denotes a Markov kernel, also known as a backward kernel, which transitions backward from $\boldsymbol{\theta}_{t+1}$ to $\boldsymbol{\theta}_{t}$. Weighted particles $\left\{w_{t}^{\left[m\right]},\boldsymbol{\theta}_{t}^{\left[m\right]}\right\}_{m=1}^{M}$ are used to approximate $\tilde{\pi}_{t}$. These particles are moved via a Markov kernel $\mathcal{K}_{t}\left(\cdot,\cdot\right)$, termed a forward kernel. Let $\gamma_{t}\left(\boldsymbol{\theta}\right)$ denote the unnormalized posterior density with $\phi_{t}$,
@@
The unnormalized weights are expressed as follows:
$W_{t}^{\left[m\right]}\propto\frac{\tilde{\pi}_{t}\left(\boldsymbol{\theta}_{t}^{\left[m\right]}\right)\prod_{s=1}^{t-1}\mathcal{L}_{s}\left(\boldsymbol{\theta}_{s+1}^{\left[m\right]},\boldsymbol{\theta}_{s}^{\left[m\right]}\right)}{\zeta_{1}\left(\boldsymbol{\theta}_{1}^{\left[m\right]}\right)\prod_{s^{\prime}=1}^{t-1}\mathcal{K}_{s^{\prime}}\left(\boldsymbol{\theta}_{s^{\prime}-1}^{\left[m\right]},\boldsymbol{\theta}_{s^{\prime}}^{\left[m\right]}\right)}.$
@@ 3 Proposed Algorithm
4 Application
4.1 Misspecified linear regression
@@
where $\boldsymbol{x}_{i}=\left(1,x_{1,i},x_{2,i},x_{3,i}\right)^{\top}$. The true data generation process, which was borrowed from Section 5.2 of [17] (Degree 2), had heteroskedastic errors depending on the first covariates:
@@
Posterior simulations were conducted using a Gibbs sampler. The unknown parameters $\boldsymbol{\theta}=\left(\boldsymbol{\beta}^{\top},\sigma^{2}\right)^{\top}$ were inferred using a normal prior, $\boldsymbol{\beta}\sim\mathcal{N}\left(\boldsymbol{0}_{4},\varsigma^{2}\boldsymbol{I}_{4}\right)$, and an inverse gamma prior, $\sigma^{2}\sim\mathcal{IG}\left(\varrho_{1},\varrho_{2}\right)$, where $\varsigma^{2}$, $\varrho_{1}$, and $\varrho_{2}$ are hyperparameters; $\mathcal{IG}\left(a,b\right)$ denotes an inverse gamma distribution with shape parameter $a$ and rate parameter $b$. The hyperparameters for $\sigma^{2}$ were set as $\left(\varrho_{1},\varrho_{2}\right)=\left(1,0.025\right)$. While [17] specified $\varsigma^{2}=\sigma^{2}$, $\boldsymbol{\beta}\sim\mathcal{N}\left(\boldsymbol{0}_{4},\sigma^{2}\boldsymbol{I}_{4}\right)$, our prior of $\boldsymbol{\beta}$ was independent of $\sigma^{2}$, $\varsigma^{2}=100$. Consequently, the resulting posterior was more general and computationally expensive because the inverse matrix had to be computed repeatedly.
@@
We applied the two algorithms to 200 synthetic datasets. As shown in Table 1, both algorithms effectively achieved the target credibility level. Figure 4 compares the computational costs of the algorithms. The panels in the left column report the wall clock time in seconds. The median computation for the GPC-WP was approximately 25-30% faster than that of the GPC-SA. The improvement in the computational efficiency was due to the smaller number of iterations, as shown in the panels in the right column. While the GPC-SA needed approximately nine iterations, the GPC-WP converged within two or three iterations in most cases. The reason why the GPC-WP was not as fast as the GPC-SA in terms of the number of iterations was that it took more time to obtain the next learning rates. Nevertheless, this additional computational cost was outweighed by the gain from reducing the number of iterations.
Table 1: Coverage probability
@@
4.2 Support vector machine
We applied the proposed algorithm to estimate a support vector machine with real data, following Section 5 of [16]. The outcome is binary, $y_{i}\in\left\{-1,1\right\}$, and the feature vector is composed of an intercept and $K-1$ covariates, $\boldsymbol{x}_{i}=\left(1,x_{1,i},...,x_{K-1,i}\right)^{\top}$. Define $\boldsymbol{y}=\left(y_{1},...,y_{N}\right)^{\top}$ and $\boldsymbol{X}=\left(\boldsymbol{x}_{1},...,\boldsymbol{x}_{N}\right)^{\top}$. Estimating the support vector classifier is represented as the problem of minimizing the following empirical risk function:
@@ 4.2 Support vector machine
where $\sigma_{k}$ denotes the standard deviation of the $k$th predictor $x_{k,1},...,x_{k,N}$ with $\sigma_{1}=1$ and $\nu\left(>0\right)$ is a hyperparameter. We chose $\nu=10$. This model was inferred for the South African Heart Disease dataset as in Section 4.4.2 of [25], with $N=462$ and $K=8$. For the posterior simulation, we modified a Gibbs sampler of [26] by incorporating $\eta$.
The two algorithms were executed 50 times using different random seeds. For every execution, both algorithms reached $\eta\approx 0.09$, which is in agreement with [16]. Figure 5 displays the computation time and number of iterations. The median computation time for the GPC-SA was 201.7 seconds, while that for the GPC-WP was 124.4 seconds (38.3% faster). The median required number of iterations for the GPC-SA was 7, while that for the GPC-WP was 4.
@@
5 Conclusion
In this paper, we explored two robust Bayesian methods: generalized Bayesian inference and Gibbs posterior inference. These methods offer robustness against model misspecifications by introducing learning rate parameters and formulating inferential problems using generic loss functions. Central to these approaches is the calibration of the posterior distribution spread through the selection of an appropriate learning rate parameter. Building upon the GPC algorithm proposed by [16], we introduced a more efficient version inspired by SMC samplers. Our proposed algorithm evaluates the coverage probability with a different learning rate without the need for repeated posterior simulations, substantially reducing the computational costs while obtaining the desired values within a few iterations. Through synthetic and real data applications, we demonstrated the efficacy of our proposed algorithm. By providing a computationally efficient solution for learning rate selection, our work contributes to advancing robust Bayesian methods in statistical learning, facilitating their practical applicability in complex modeling scenarios.
In future work, further exploration of the algorithm’s performance under diverse model structures and datasets could provide additional insights and enhancements. In addition, [27] recently proposed an approach that delivers a well-calibrated generalized posterior without a learning rate parameter. A thorough comparison with our algorithm is beyond the scope of this paper, and further exploration is left for future work.

### v3

@@ Abstract
In the realm of statistical learning, the increasing volume of accessible data and increasing model complexity necessitate robust methodologies. This paper explores two branches of robust Bayesian methods in response to this trend. The first is generalized Bayesian inference, which introduces a learning rate parameter to enhance robustness against model misspecifications. The second is Gibbs posterior inference, which formulates inferential problems using generic loss functions rather than probabilistic models. In such approaches, it is necessary to calibrate the spread of the posterior distribution by selecting a learning rate parameter. The study aims to enhance the generalized posterior calibration (GPC) algorithm proposed by [1]. Their algorithm chooses the learning rate to achieve the nominal frequentist coverage probability, but it is computationally intensive because it requires repeated posterior simulations for bootstrap samples. We propose a more efficient version of the GPC inspired by sequential Monte Carlo (SMC) samplers. A target distribution with a different learning rate is evaluated without posterior simulation as in the reweighting step in SMC sampling. Thus, the proposed algorithm can reach the desirable value within a few iterations. This improvement substantially reduces the computational cost of the GPC. Its efficacy is demonstrated through synthetic and real data applications.
I Introduction
@@
$\pi\left(\boldsymbol{\theta}\right)\propto p\left(\mathcal{D}|\boldsymbol{\theta}\right)p\left(\boldsymbol{\theta}\right),$
where $\propto$ denotes a proportional relationship. In contrast, generalized Bayesian inference introduces a generalized posterior by incorporating a learning rate $\eta\left(>0\right)$ (also termed a scaling parameter) alongside the likelihood, as follows:
@@
This paper proposes an improvement in the GPC that requires fewer of iterations. This algorithm is inspired by sequential Monte Carlo (SMC) samplers [18], exploiting the similarity between the learning rate in the GPC and an inverse temperature in SMC sampling. The target distribution with the current value of $\eta$ is approximated using weighted particles, as in SMC sampling. The particle approximation of the target distribution with a new value of $\eta$ is obtained without posterior simulation as in the reweighting step in SMC sampling. At least in theory, a desirable value of $\eta$ is reached in a single step, although few iterations are required in practice due to the particle degeneracy problem.
Two recently proposed approaches are relevant to this paper; however, they are limited in the dimension of the problem they can handle. First, [19] proposed an algorithm that replaces an MCMC sampler with an SMC sampler. The algorithm proposed in this paper is different, in that it uses an MCMC sampler for posterior simulation and improves on the optimization procedure. Whereas SMC samplers are only applicable to relatively low-dimensional cases (dimension less than 10); however, some MCMC samplers are feasible for high-dimensional cases. Second, [20] proposed an approach that delivers a well-calibrated generalized posterior without a learning rate parameter, which is theoretically different from the posterior considered in GPC. The construction of quasi-posteriors [20] is similar to that of the quasi-posterior based on generalized method of moments criterion (e.g., [21]). Conducting Monte Carlo simulations on this type of quasi-posterior is known to be very difficult [22] requiring that the posterior distribution be evaluated each time and the inverse and determinant of a matrix of the same dimension as the number of unknown parameters be computed. Therefore, unlike the approach proposed in this paper, efficient posterior simulators such as Gibbs samplers are not available for their framework.
@@
II Background
II-A Generalized posterior calibration
@@
where $\mathbb{I}\left\{\cdot\right\}$ denotes the indicator function. The learning rate $\eta$ is selected by solving $\hat{c}_{\alpha}\left(\eta|\mathbb{P}_{N}\right)=1-\alpha$ via a stochastic approximation [23]. At the $s$th iteration, a single step of stochastic approximation recursion is applied:
@@
where $\left\{\varsigma_{l}\right\}$ is a nonincreasing sequence such that $\sum_{l}\varsigma_{l}=\infty$ and $\sum_{l}\varsigma_{l}^{2}<\infty$. While [1] specified $\varsigma_{s}=s^{-0.51}$, we adopt a variant of Keston’s [24] rule (Figure 2), $\varsigma_{s}\left(l\right)=l^{-0.51}$, where an auxiliary variable $l$ increases by one only when there is a directional change in the trajectory of $\eta_{s}$ and $\hat{c}_{\alpha}\left(\eta_{s}|\mathbb{P}_{N}\right)<1$. This adaptation significantly reduces the convergence time.
@@
II-B Sequential Monte Carlo sampler
The following describles sequential Monte Carlo (SMC) samplers [18]. These algorithms repeatedly utilize importance sampling to generate $M$ weighted particles $\left\{w^{\left[m\right]},\boldsymbol{\theta}^{\left[m\right]}\right\}_{m=1}^{M}$ that approximate a sequence of synthetic intermediate distributions $\left\{\pi_{t}\right\}_{t=0}^{T}$, where $\boldsymbol{\theta}^{\left[m\right]}$ denotes the $m$th particle and $w^{\left[m\right]}$ denotes the corresponding weight. Initially, the distribution $\pi_{0}$ corresponds to the prior, $\pi_{0}\left(\boldsymbol{\theta}\right)=p\left(\boldsymbol{\theta}\right)$, while the terminal distribution $\pi_{T}$ is the target distribution, i.e., the posterior, $\pi_{T}\left(\boldsymbol{\theta}\right)=p\left(\mathcal{D}|\boldsymbol{\theta}\right)p\left(\boldsymbol{\theta}\right)$. Each intermediate distribution is defined as a likelihood-tempered posterior:
@@ II-B Sequential Monte Carlo sampler
where $\left\{\phi_{t}\right\}_{t=1}^{T}$ is an increasing sequence with $\phi_{0}=0$ and $\phi_{T}=1$. The parameter $\phi_{t}$ can be interpreted as the inverse temperature. We denote a sequence of auxiliary distributions as $\left\{\tilde{\pi}_{t}\right\}_{t=1}^{T}$ with
$\tilde{\pi}_{t}\left(\boldsymbol{\theta}_{1:t}\right)=\pi_{t}\left(\boldsymbol{\theta}_{t}\right)\prod_{s=1}^{t-1}\mathcal{L}_{s}\left(\boldsymbol{\theta}_{s+1},\boldsymbol{\theta}_{s}\right),$
where $\mathcal{L}_{t}\left(\cdot,\cdot\right)$ denotes a Markov kernel, also known as a backward kernel, which transitions backward from $\boldsymbol{\theta}_{t+1}$ to $\boldsymbol{\theta}_{t}$. Weighted particles $\left\{w_{t}^{\left[m\right]},\boldsymbol{\theta}_{t}^{\left[m\right]}\right\}_{m=1}^{M}$ are used to approximate $\tilde{\pi}_{t}$. These particles are moved via a Markov kernel $\mathcal{K}_{t}\left(\cdot,\cdot\right)$, also termed a forward kernel. Let $\gamma_{t}\left(\boldsymbol{\theta}\right)$ denote the unnormalized posterior density with $\phi_{t}$,
@@
The sampling distribution of $\boldsymbol{\theta}_{1:t}^{\left[m\right]}=\left\{\boldsymbol{\theta}_{1}^{\left[m\right]},...,\boldsymbol{\theta}_{t}^{\left[m\right]}\right\}$ is represented as:
$\zeta_{t}\left(\boldsymbol{\theta}_{1:t}^{\left[m\right]}\right)=\zeta_{1}\left(\boldsymbol{\theta}_{1}^{\left[m\right]}\right)\prod_{s^{\prime}=2}^{t}\mathcal{K}_{s^{\prime}}\left(\boldsymbol{\theta}_{s^{\prime}-1}^{\left[m\right]},\boldsymbol{\theta}_{s^{\prime}}^{\left[m\right]}\right).$
Importance sampling is applied to correct the discrepancy between $\tilde{\pi}_{t}$ and $\zeta_{t}$. The unnormalized weights are expressed as follows:
$W_{t}^{\left[m\right]}\propto\frac{\tilde{\pi}_{t}\left(\boldsymbol{\theta}_{1:t}^{\left[m\right]}\right)}{\zeta_{t}\left(\boldsymbol{\theta}_{1:t}^{\left[m\right]}\right)}=\frac{\tilde{\pi}_{t}\left(\boldsymbol{\theta}_{t}^{\left[m\right]}\right)\prod_{s=1}^{t-1}\mathcal{L}_{s}\left(\boldsymbol{\theta}_{s+1}^{\left[m\right]},\boldsymbol{\theta}_{s}^{\left[m\right]}\right)}{\zeta_{1}\left(\boldsymbol{\theta}_{1}^{\left[m\right]}\right)\prod_{s^{\prime}=2}^{t}\mathcal{K}_{s^{\prime}}\left(\boldsymbol{\theta}_{s^{\prime}-1}^{\left[m\right]},\boldsymbol{\theta}_{s^{\prime}}^{\left[m\right]}\right)}.$
@@ III Proposed Algorithm
IV Application
IV-A Misspecified linear regression
@@
where $y_{i}$ denotes a response variable; $\boldsymbol{x}_{i}=\left(1,x_{1,i},x_{2,i},x_{3,i}\right)^{\top}$ denotes a vector of covariates; $\boldsymbol{\beta}$ is the corresponding coefficient vector; $\varepsilon_{i}$ denotes an error term with variance $\sigma^{2}$; $\mathcal{N}\left(a,b^{2}\right)$ denotes a normal distribution with mean $a$ and variance $b^{2}$. The true data generation process, which was borrowed from Section 5.2 of [17] (Degree 2), had heteroskedastic errors depending on $x_{1,i}$:
@@
Posterior simulations were conducted using a Gibbs sampler. The unknown parameters $\boldsymbol{\theta}=\left(\boldsymbol{\beta}^{\top},\sigma^{2}\right)^{\top}$ were inferred using a normal prior, $\boldsymbol{\beta}\sim\mathcal{N}\left(\boldsymbol{0}_{4},\varsigma^{2}\boldsymbol{I}_{4}\right)$, and an inverse gamma prior, $\sigma^{2}\sim\mathcal{IG}\left(\varrho_{1},\varrho_{2}\right)$, where $\boldsymbol{0}_{A}$ denotes the $A$-dimensional vector of zeros, $\boldsymbol{I}_{A}$ denotes the $A$-dimensional identity matrix, $\varsigma^{2}$, $\varrho_{1}$, and $\varrho_{2}$ are hyperparameters; $\mathcal{IG}\left(a,b\right)$ denotes an inverse gamma distribution with shape parameter $a$ and rate parameter $b$. The hyperparameters for $\sigma^{2}$ were set as $\left(\varrho_{1},\varrho_{2}\right)=\left(1,0.025\right)$. While [17] specified $\varsigma^{2}=\sigma^{2}$, $\beta_{i}\sim\mathcal{N}\left(0,\sigma^{2}\right)$, our prior of $\boldsymbol{\beta}$ was independent of $\sigma^{2}$, $\varsigma^{2}=100$. Consequently, the resulting posterior was more general and computationally expensive because the inverse matrix had to be computed repeatedly.
@@
We applied the two algorithms to 200 synthetic datasets. 3,000 posterior draws were simulated and the last 2,000 draws were used for posterior analysis. As shown in Table I, both algorithms effectively achieved the target credibility level. Table II and Figure 4 compares the computational costs of the algorithms. The panels in the left column report the wall clock time in seconds. The median computation for the GPC-WP was approximately 25-30% faster than that of the GPC-SA. The improvement in the computational efficiency was due to the smaller number of iterations, as shown in the right two columns of Table III and the the panels in the right column of Figure 4. While the GPC-SA needed approximately nine iterations, the GPC-WP converged within two or three iterations in most cases. The reason why the GPC-WP was not as fast as the GPC-SA in terms of the number of iterations was that it took more time to obtain the next learning rates. Nevertheless, this additional computational cost was outweighed by the gain from reducing the number of iterations.
Table I: Coverage probability
@@
Table II: Comparison of the computational cost
$N$ Computation time (sec.) Number of iteration
GPC-SA GPC-WP GPC-SA GPC-WP
100 12.3 8.8 9 3
(9.7, 13.8) (7.9, 10.0) (7,10) (2,3)
200 12.4 9.2 9 3
(11.0, 13.8) (8.4, 10.3) (8,10) (2,3)
400 12.6 9.6 9 3
(11.2, 14.0) (8.6, 10.4) (8,10) (2,3)
Note: The median value is computed based on 200 synthetic datasets. The numbers in parenthesis denote the interquartile range.
@@
IV-B Support vector machine
We applied the proposed algorithm to estimate a support vector machine with real data, following Section 5 of [1]. The outcome is binary, $y_{i}\in\left\{-1,1\right\}$, and the feature vector is composed of an intercept and $K-1$ covariates, $\boldsymbol{x}_{i}=\left(1,x_{1,i},...,x_{K-1,i}\right)^{\top}$. Estimating the support vector classifier is represented as the problem of minimizing the following empirical risk function:
@@ IV-B Support vector machine
where $\sigma_{k}$ denotes the standard deviation of the $k$th predictor $x_{k,1},...,x_{k,N}$ with $\sigma_{1}=1$, and $\nu\left(>0\right)$ is a hyperparameter. We chose $\nu=10$. This model was inferred for the South African Heart Disease dataset as in Section 4.4.2 of [28], with $N=462$ and $K=8$. For the posterior simulation, we modified a Gibbs sampler of [29] by incorporating $\eta$.
The two algorithms were executed 50 times using different random seeds. We generated total 11,000 draws, discarding the first 1,000 draws as warmup. For every execution, both algorithms reached $\eta\approx 0.09$, which is in agreement with [1]. Table III and Figure 5 report the computation time and number of iterations. The median computation time for the GPC-SA was 201.7 seconds, while that for the GPC-WP was 124.4 seconds (38.3% faster). The median required number of iterations for the GPC-SA was 7, while that for the GPC-WP was 4.
Table III: Comparison of the computational cost
Computation time (sec.) Number of iteration
GPC-SA GPC-WP GPC-SA GPC-WP
201.7 124.4 7 4
(144.6, 231.4) (93.2, 183.9) (5, 8) (3, 5)
Note: The median value is computed based on 200 synthetic datasets. The numbers in parenthesis denote the interquartile range.
@@
V Conclusion
In this paper, we explored two robust Bayesian methods: generalized Bayesian inference and Gibbs posterior inference. These methods offer robustness against model misspecifications by introducing learning rate parameters and formulating inferential problems using generic loss functions. Central to these approaches is the calibration of the posterior distribution spread through the selection of an appropriate learning rate parameter. Building upon the GPC algorithm proposed by [1], we introduced a more efficient version inspired by SMC samplers. Our proposed algorithm evaluates the coverage probability with a different learning rate without the need for repeated posterior simulations, substantially reducing the computational costs while obtaining the desirable values within a few iterations. Through synthetic and real data applications, we demonstrated the efficacy of our proposed algorithm. By providing a computationally efficient solution for learning rate selection, our work contributes to advancing robust Bayesian methods in statistical learning, facilitating their practical applicability in complex modeling scenarios.